# Harnessing Large Language Models as Post-hoc Correctors

### v1

@@ Abstract
As Machine Learning (ML) models grow in size and demand higher-quality training data, the expenses associated with re-training and fine-tuning these models are escalating rapidly. Inspired by recent impressive achievements of Large Language Models (LLMs) in different fields, this paper delves into the question: can LLMs efficiently improve an ML’s performance at a minimal cost? We show that, through our proposed training-free framework LlmCorr, an LLM can work as a post-hoc corrector to propose corrections for the predictions of an arbitrary ML model. In particular, we form a contextual knowledge database by incorporating the dataset’s label information and the ML model’s predictions on the validation dataset. Leveraging the in-context learning capability of LLMs, we ask the LLM to summarise the instances in which the ML model makes mistakes and the correlation between primary predictions and true labels. Following this, the LLM can transfer its acquired knowledge to suggest corrections for the ML model’s predictions. Our experimental results on the challenging molecular predictions show that LlmCorr improves the performance of a number of models by up to $39\%$
@@ 1 Introduction
In recent decades, Machine Learning (ML) models have become increasingly prevalent in various real-world applications Butler et al. (2018); Dixon et al. (2020); Zhong et al. (2023). As ML models grow in size and demand higher-quality training data, the expenses associated with pre-training and fine-tuning these models to achieve superior performances are escalating rapidly Devlin et al. (2018); He et al. (2021). Hence, there is an urgent need to develop effective, lightweight and practical solutions for users to improve their ML model’s predictions.
@@
To validate the effectiveness of LlmCorr, we deploy it in the face of structured molecule graph data within the domain of biology, e.g., predicting the functionality of molecules. Through extensive experiments conducted on six real-world benchmark datasets covering diverse subjects, we empirically demonstrate that LlmCorr significantly elevates the quality of predictions across diverse ML models, achieving notable improvements, up to 39%. Furthermore, we conduct comprehensive follow-up ablation studies and analyses to validate the efficacy of LlmCorr’s designs and elucidate the impact of key factors.
@@ 3 Preliminary and Problem
This paper aims to leverage LLMs as post-hoc correctors to enhance predictions made by an arbitrary ML model. Specifically, we focus on addressing challenging prediction tasks on structured molecule graph data within the field of biology. For instance, consider the supervised molecule property prediction task, where molecules can be represented using various formats such as SMILES string Weininger (1988) and geometry structures Zhang et al. (2024) (as shown in Figure 8). However, a notable limitation of existing LLMs is their reliance on unstructured text, rendering them unable to incorporate essential geometry structures as input Li et al. (2023); Guo et al. (2023). To overcome this limitation, Fatemi et al. (2023) propose encoding the graph structure into text descriptions. In this paper, as depicted in Figure 8, we extend this concept by encoding both the molecule’s atom features and graph structure into textual descriptions.
@@ 4.1 Contextual Knowledge Database Construction
After completing the training of the ML model ($f_{ML}$) on the training set $\mathcal{M}_{train}$, we gather comprehensive data from both the training set $\mathcal{M}_{train}$ and the validation set $\mathcal{M}_{val}$, along with the primary predictions $\hat{\mathcal{Y}}_{val}$ made by $f_{ML}$ on the validation set. Subsequently, we construct a contextual knowledge database $\mathcal{D}$ based on the collected data. Notably, the knowledge database $\mathcal{D}$ not only contains information regarding the original training and validation datasets but also provides insights into the types of molecules for which the ML models generate accurate or inaccurate predictions. Additionally, it captures the relationship between the initial prediction $\hat{\mathcal{Y}}_{val}$ and the true label $\mathcal{Y}_{val}$. We anticipate that this essential contextual knowledge will empower the LLM ($f_{LLM}$) to refine the predictions $\hat{\mathcal{Y}}_{test}$ made by $f_{ML}$ on the test dataset $\mathcal{M}_{test}$.
@@ 4.3 LlmCorr Prompt Engineering
    Contextual knowledge from the training dataset: Includes SMILES string, molecule atom feature and structure descriptions, and molecule label information of the training dataset.
@@
    Contextual knowledge from the validation dataset: Includes SMILES string, molecule atom feature and structure descriptions, molecule label information and the ML model’s predictions of the validation dataset. This equips the LLM with insights into the ML model’s error patterns, enhancing its ability to refine predictions on the test dataset.
@@
Extensibility. LlmCorr is designed as a post-hoc corrector framework that leverages the ICL capability of LLMs. As illustrated in Figure 2, LlmCorr is adaptable to an arbitrary ML model, showcasing its remarkable extensibility. With the trend of ML models increasing in size and requiring higher-quality training data, the costs associated with enhancing ML models by re-training and fine-tuning are rapidly escalating. LlmCorr emerges as a promising, general-purpose practical solution in the LLM era. LlmCorr is summarised in Algorithm 1 in Appendix B.
@@ 5 Are LLMs Post-hoc Correctors?
In this section, we evaluate the effectiveness of LLMs as post-hoc correctors. Our experimental analysis focuses on the challenging structured molecular graph property prediction tasks.
Dataset. We consider six widely used molecule datasets from the OGB benchmark Hu et al. (2021b), including ogbg-molbace, ogbg-molbbbp, ogbg-molhiv, ogbg-molesol, ogbg-molfreesolv and ogbg-mollipo. Detailed descriptions are summarised in Appendix C.
ML Models. To investigate whether LlmCorr can effectively improve predictions across various types of ML models. We consider ML models of three different categories: (1) Language Model (LM) that only takes text information as inputs, i.e., DeBERTa He et al. (2021). (2) Graph Neural Networks (GNNs) that capture the molecule’s geometry structure information. We consider two classic GNN variants, i.e., GCN Kipf and Welling (2017) and GIN Xu et al. (2019), and two SOTA GNN variants collected from the OGB leaderboards Hu et al. (2021b), i.e., HIG and PAS Wei et al. (2021). (3) And we consider one recently released hybrid framework, TAPE He et al. (2023), in our experiments. TAPE leverages LM and LLMs to capture textual information as features, which can be used to boost GNN performance. The implementation details are discussed in Appendix D.
LLMs. In this work, we are interested in where the LLM’s parameters are fixed, and the system is available for users in a black box setup where the LLM only consumes and produces text. We believe this setting to be particularly valuable as most users would practically have access to LLMs. In this case, we consider GPT-3.5 and GPT-4 Achiam et al. (2023) as LLMs in this work, and GPT-3.5 is the major LLM for most experiments. All responses are obtained by calling their official APIs. Because the generated descriptions following Fatemi et al. (2023) have tons of tokens, easily over the LLM’s input token constraints, hence we do not include descriptions in the LlmCorr prompt in this study.
@@ 5.1 Main Results
Observation 4: GPT-4 underperforms compared to GPT-3.5. Table 2 displays the molecule graph property prediction performance and execution time for three datasets, comparing GPT-3.5 and GPT-4. Interestingly, we find that despite its larger training data and more complex fine-tuning process, GPT-4 exhibits inferior performance compared to GPT-3.5 in this study. We hypothesise that this discrepancy may be attributed to interventions such as reinforcement learning through human feedback.
@@
Observation 5: LLMs exhibit limited competitiveness as predictors. Given LlmCorr’s remarkable performance as correctors, another intriguing question arises: can LLM generate accurate predictions directly? To investigate, we conduct additional experiments where the LLM is tasked with directly predicting the molecule’s property. For detailed findings, please refer to Appendix E due to space constraints. As shown in the results of Table 3, LLMs do not demonstrate competitive performance as predictors. This observation reinforces the efficacy of LlmCorr, which leverages LLMs as post-hoc correctors.
@@ 5.2 Ablation Study
Table 4: Ablation study of LlmCorr on ogbg-molbace and ogbg-molesol with variants of contextual knowledge retrieval.
ogbg-molbace ogbg-molesol
ROC-AUC $\uparrow$ RMSE $\downarrow$
@@ 7 Limitations and Ethic Statement
Limitations. While LlmCorr demonstrates simplicity and effectiveness in improving the predictions of an arbitrary ML model, our verification was confined to structured molecular graph property prediction tasks. Further extensive empirical investigations across diverse domains are warranted to establish its generalisability. Additionally, considering the purported enhanced ICL capabilities of GPT-4 on various benchmark tasks OpenAI (2023), it is noteworthy that our findings (as discussed in Section 5.1 and illustrated in Table 2) reveal GPT-4’s underperformance compared to the GPT-3.5 model. This discrepancy merits further exploration to elucidate the underlying reasons. Moreover, while LlmCorr’s prompt template accommodates the insertion of molecule atom features and geometry structure descriptions, limitations stemming from the LLM’s input token constraints prevented their inclusion in the prompt. Lastly, while our approach incorporates contextual knowledge into the prompt, its utility is constrained by several factors, including limited flexibility. For example, further leveraging different techniques, e.g., RAG Lewis et al. (2020), to involve more contextual knowledge into the LLM is also a fruitful direction. Further enhancements in this regard are warranted to maximise LlmCorr’s effectiveness.
@@ Appendix A Illustration of Molecule Representations
Appendix B Algorithm
@@ 7       Query the LLM and contain the refined prediction $\tilde{y}_{u}=f_{LLM}(P_{u})$
Appendix C Dataset Description
@@
Appendix D Implementation
Implementation. We implement ML predictive models following their available official implementations. For instance, we adopt the available code of variant GNN models on the OGB benchmark leaderboards, e.g., GCN [^1] https://github.com/snap-stanford/ogb/tree/master/examples/graphproppred/mol, GIN [^2] https://github.com/snap-stanford/ogb/tree/master/examples/graphproppred/mol, HIG  [^3] https://github.com/TencentYoutuResearch/HIG-GraphClassification and PAS  [^4] https://github.com/LARS-research/PAS-OGB. About DeBERTa, we adopt its official implementation  [^5] https://huggingface.co/microsoft/deberta-v3-base and incorporate it within the pipeline of TAPE  [^6] https://github.com/XiaoxinHe/TAPE. For the LLMs, we simply call the API provided by OpenAI with default hyper-parameter settings. We empirically tried with some combinations of recommended important hyper-parameters, e.g., temperature and top_P, yet did not observe significant improvement. To realise the embedding-based information retrieval for LlmCorr, we adopt two capable embedding models ($f_{Emb}$) provided by OpenAI  [^7] https://platform.openai.com/docs/models/embeddings, e.g., text-embedding-3-large and text-embedding-3-small. In this work, we mainly adopt text-embedding-3-large for better empirical performance. We perform careful discussions about the influence of different variants in Section 5.2.
Appendix E Are LLMs Predictors?
@@
Figure 9: Zero-shot prompt templates.
Figure 10: Few-shot prompt template. Multiple contextual knowledge can be included by expanding the template.
E.1 Predictor Prompt Engineering
Zero-shot Prompting. The first set of prompts (IP, IE) simply provides the LLM with molecule and task descriptions and asks it to generate the desired output with a desired format without any prior training or knowledge on the task, as illustrated in Figure 9. The only guidance we provide to the LLM is instruction, which tells about a little background context. Particularly, IP only asks the LLM to provide predictions, while IE further asks for explanations, which may ask the LLM to clarify the thought process in explanation generation and provide helpful evidence to help users understand the given prediction. In addition, if we fill out the description of IP and IE, which derives IPD and IPD prompts.
Few-shot Prompting. The second kind of prompt (FS) that we propose provides the LLM with a small number of examples of the task, along with the desired outputs Brown et al. (2020). The model then learns from these examples to perform the task on new inputs. This approach can be categorised as a simple in-context learning (ICL) technique, An example prompt template is shown in Figure 10. FS-$k$ indicates $k$ contextual knowledge instances are included in the prompt. In this work, we do not discuss the FSD prompts since the generated descriptions have tons of tokens, which will easily go over the LLM’s input constraints.
We note there are also some popular recent ICL techniques, e.g., Chain-of-thought (CoT) Wei et al. (2022b), Tree-of-thought (ToT) Yao et al. (2023), Graph-of-thought (GoT) Besta et al. (2023) and Retrieval Augmented Generation (RaG) Lewis et al. (2020), which are theoretically available to support complicated tasks and include large knowledge context. However, our initial experiments showed that methods, e.g., CoT, ToT and GoT, perform much worse for molecule property prediction tasks due to the significant difficulties in designing proper chain thoughts without solid expertise. RaG implementations that we tested are unstable and slow with query, and they fall short of the relatively simpler FS’s performance. We argue it is caused by the unqualified information retrieval system, and we will discuss it in the future work discussion section.
E.2 Results - LLMs work as Predictors

### v2

@@ Abstract
As Machine Learning (ML) models grow in size and demand higher-quality training data, the expenses associated with re-training and fine-tuning these models are escalating rapidly. Inspired by recent impressive achievements of Large Language Models (LLMs) in different fields, this paper delves into the question: can LLMs efficiently improve an ML’s performance at a minimal cost? We show that, through our proposed training-free framework LlmCorr, an LLM can work as a post-hoc corrector to propose corrections for the predictions of an arbitrary ML model. In particular, we form a contextual knowledge database by incorporating the dataset’s label information and the ML model’s predictions on the validation dataset. Leveraging the in-context learning capability of LLMs, we ask the LLM to summarise the instances in which the ML model makes mistakes and the correlation between primary predictions and true labels. Following this, the LLM can transfer its acquired knowledge to suggest corrections for the ML model’s predictions. Our experimental results on text analysis and the challenging molecular predictions show that LlmCorr improves the performance of a number of models by up to $39\%$  [^1] The code and models are available at https://github.com/zhiqiangzhongddu/LLMCorr..
[…]
[backgroundcolor=gray!20, linecolor=red!60!black, linewidth=2pt, topline=false, rightline=false, skipabove=10pt, skipbelow=10pt, leftline=false]regboxBox
@@ 1 Introduction
In recent decades, Machine Learning (ML) models have become increasingly prevalent in various real-world applications Dixon et al. (2020); Zhong et al. (2023). As ML models grow in size and demand higher-quality training data, the expenses associated with pre-training and fine-tuning these models to achieve superior performances are escalating rapidly Devlin et al. (2018); He et al. (2021). Hence, there is an urgent need to develop effective, lightweight and practical solutions for users to improve their ML model’s predictions.
@@
To validate the effectiveness of LlmCorr, we mainly deploy it in the face of structured molecule graph data within the domain of biology, e.g., predicting the functionality of molecules. Through extensive experiments conducted on six real-world benchmark datasets covering diverse subjects, we empirically demonstrate that LlmCorr significantly elevates the quality of predictions across diverse ML models, achieving notable improvements, up to 39%. Besides, we demonstrate LlmCorr’s effectiveness on general text analysis tasks. Furthermore, we conduct comprehensive follow-up ablation studies and analyses to validate the efficacy of LlmCorr’s designs and elucidate the impact of key factors.
@@ 3 Preliminary and Problem
This paper aims to leverage LLMs as post-hoc correctors to enhance predictions made by an arbitrary ML model. Specifically, we showcase our framework on challenging prediction tasks on structured molecule graph data in biology. We further demonstrate the generality of our framework on sentiment analysis and text classification in Section  in the Appendix. For instance, consider the supervised molecule property prediction task, where molecules can be represented using various formats such as SMILES string Weininger (1988) and geometry structures Zhang et al. (2024) (as shown in Figure 8). However, a notable limitation of existing LLMs is their reliance on unstructured text, rendering them unable to incorporate essential geometry structures as input Li et al. (2023); Guo et al. (2023). To overcome this limitation, Fatemi et al. (2023) propose encoding the graph structure into text descriptions. In this paper, as depicted in Figure 8, we extend this concept by encoding both the molecule’s atom features and graph structure into textual descriptions.
@@ 4.1 Contextual Knowledge Database Construction
After completing the training of the ML model ($f_{ML}$) on the training set $\mathcal{M}_{train}$, we collect data from both the training set $\mathcal{M}_{train}$ and the validation set $\mathcal{M}_{val}$, along with the primary predictions $\hat{\mathcal{Y}}_{val}$ made by $f_{ML}$ on the validation set. Subsequently, we construct a contextual knowledge database $\mathcal{D}$ based on the collected data. Notably, the knowledge database $\mathcal{D}$ not only contains information regarding the original training and validation datasets but also provides insights into the types of molecules for which the ML models generate accurate or inaccurate predictions. Additionally, it captures the relationship between the initial prediction $\hat{\mathcal{Y}}_{val}$ and the true label $\mathcal{Y}_{val}$. We anticipate that this essential contextual knowledge will empower the LLM ($f_{LLM}$) to refine the predictions $\hat{\mathcal{Y}}_{test}$ made by $f_{ML}$ on the test dataset $\mathcal{M}_{test}$.
@@ 4.3 LlmCorr Prompt Engineering
    Contextual knowledge from the training dataset: Includes SMILES string and molecule label information of the training dataset.
@@
    Contextual knowledge from the validation dataset: Includes SMILES string, molecule label information and the ML model’s predictions of the validation dataset. This equips the LLM with insights into the ML model’s error patterns, enhancing its ability to refine predictions on the test dataset.
@@
Extensibility. LlmCorr is designed as a post-hoc corrector framework that leverages the ICL capability of LLMs. As illustrated in Figure 2, LlmCorr is adaptable to an arbitrary ML model, showcasing its remarkable extensibility. With the trend of ML models increasing in size and requiring higher-quality training data, the costs associated with enhancing ML models by re-training and fine-tuning are rapidly escalating. LlmCorr emerges as a promising, general-purpose practical solution in the LLM era. LlmCorr is summarised in Algorithm 1 in Appendix C.
@@ 5 Are LLMs Post-hoc Correctors?
In this section, we evaluate the effectiveness of LLMs as post-hoc correctors. Our experimental analysis mainly focuses on the challenging structured molecular graph property prediction tasks. The additional experiments on general text analysis tasks are presented in Appendix G.
Dataset. We consider six widely used molecule datasets from the OGB benchmark Hu et al. (2021b), including ogbg-molbace, ogbg-molbbbp, ogbg-molhiv, ogbg-molesol, ogbg-molfreesolv and ogbg-mollipo. Detailed descriptions are summarised in Appendix D.
ML Models. To investigate whether LlmCorr can effectively improve predictions across various types of ML models. We consider ML models of three different categories: (1) Language Model (LM) that only takes text information as inputs, i.e., DeBERTa He et al. (2021). (2) Graph Neural Networks (GNNs) that capture the molecule’s geometry structure information. We consider two classic GNN variants, i.e., GCN Kipf and Welling (2017) and GIN Xu et al. (2019), and two SOTA GNN variants collected from the OGB leaderboards Hu et al. (2021b), i.e., HIG and PAS Wei et al. (2021). (3) And we consider one recently released hybrid framework, TAPE He et al. (2023), in our experiments. TAPE leverages LM and LLMs to capture textual information as features, which can be used to boost GNN performance. The implementation details are discussed in Appendix E.
LLMs. In this work, we are interested in where the LLM’s parameters are fixed, and the system is available for users in a black box setup where the LLM only consumes and produces text. We believe this setting to be particularly valuable as most users would practically have access to LLMs. In this case, we consider GPT-3.5 and GPT-4 Achiam et al. (2023) as LLMs in this work, and GPT-3.5 is the major LLM for most experiments. Besides, open-source LLMs are also of great interest since they are easier to deploy privately, so we adopt Llama2 Touvron et al. (2023) as another LLM. Because the generated descriptions following Fatemi et al. (2023) have tons of tokens, easily over the LLM’s input token constraints, hence we do not include descriptions in the LlmCorr prompt in this study. Including the description information is a promising future work to explore with LLMs of more input tokens.
@@ 5.1 Main Results
[…]
Observation 4: GPT-4 and Llama2 underperform compared to GPT-3.5. Table 2 displays the molecule graph property prediction performance and execution time for three datasets, comparing Llama2, GPT-3.5 and GPT-4. Llama2 has fewer parameters than GPT-3.5 and GPT-4, but Llama2 can still improve the predictions within the framework of LlmCorr on some datasets. Interestingly, we find that despite its larger training data and more complex fine-tuning process, GPT-4 exhibits inferior performance compared to GPT-3.5 in this study. We hypothesise that this discrepancy may be attributed to interventions such as reinforcement learning through human feedback.
@@
Observation 5: LLMs exhibit limited competitiveness as predictors. Given LlmCorr’s remarkable performance as correctors, another intriguing question arises: can LLM generate accurate predictions directly? To investigate, we conduct additional experiments where the LLM is tasked with directly predicting the molecule’s property. For detailed findings, please refer to Appendix F due to space constraints. As shown in the results of Table 3, LLMs do not demonstrate competitive performance as predictors. This observation reinforces the efficacy of LlmCorr, which leverages LLMs as post-hoc correctors.
@@ 5.2 Ablation Study
Table 4: Ablation study of LlmCorr on ogbg-molbace, ogbg-molesol and ogbg-mollipo with variants of contextual knowledge retrieval.
ogbg-molbace ogbg-molesol ogbg-mollipo
ROC-AUC $\uparrow$ RMSE $\downarrow$ RMSE $\downarrow$
@@ 7 Limitations and Ethic Statement
Limitations. While LlmCorr demonstrates simplicity and effectiveness in improving the predictions of an arbitrary ML model, our verification was mainly confined to structured molecular graph property prediction tasks and several text analysis tasks. Further extensive empirical investigations across diverse domains are warranted to establish its generalisability. Additionally, considering the purported enhanced ICL capabilities of GPT-4 on various benchmark tasks OpenAI (2023), it is noteworthy that our findings (as discussed in Section 5.1 and illustrated in Table 2) reveal GPT-4’s underperformance compared to the GPT-3.5 model. This discrepancy merits further exploration to elucidate the underlying reasons. Moreover, while LlmCorr’s prompt template could accommodate the insertion of molecule atom features and geometry structure descriptions, similar to prompt templates shown in Appendix F, limitations stemming from the LLM’s input token constraints prevented their inclusion in the prompt in this paper. We believe that with the rapid development of LLMs, some LLMs that allow much longer inputs will soon be available. It would be interesting to investigate LlmCorr’s effectiveness while including descriptions in the prompt. Lastly, while our approach incorporates contextual knowledge into the prompt, its utility is constrained by several factors, including limited flexibility. For example, further leveraging different techniques, e.g., RAG Lewis et al. (2020) to involve more contextual knowledge into the LLM and advanced contextual knowledge retrieval Zhang et al. (2022) are also fruitful directions. Further enhancements in this regard are warranted to maximise LlmCorr’s effectiveness.
@@ Appendix A Illustration of Molecule Representations
Appendix B LlmCorr Prompt Templates for Regression Tasks
Due to the page space limit, we only provide LlmCorr prompt templates in Section 4. This section illustrates our LlmCorr prompt templates for regression tasks. Particularly, the proposed LlmCorr prompt template for regression tasks is illustrated in Figure 9, which contains similar components as LlmCorr prompt template for classification tasks as discussed in Section 4.3. The self-construction prompt for regression tasks is illustrated in Figure 10.
Figure 9: LlmCorr prompt template for regression tasks. Multiple contextual knowledge from training and validation datasets can be included by expanding the template.
Figure 10: Self-correction prompt template for regression tasks.
Appendix C Algorithm
@@ 7       Query the LLM and contain the refined prediction $\tilde{y}_{u}=f_{LLM}(P_{u})$
Appendix D Dataset Description
@@
Appendix E Implementation
Implementation. We implement ML predictive models following their available official implementations. For instance, we adopt the available code of variant GNN models on the OGB benchmark leaderboards, e.g., GCN [^2] https://github.com/snap-stanford/ogb/tree/master/examples/graphproppred/mol, GIN [^3] https://github.com/snap-stanford/ogb/tree/master/examples/graphproppred/mol, HIG  [^4] https://github.com/TencentYoutuResearch/HIG-GraphClassification and PAS  [^5] https://github.com/LARS-research/PAS-OGB. About DeBERTa, we adopt its official implementation  [^6] https://huggingface.co/microsoft/deberta-v3-base and incorporate it within the pipeline of TAPE  [^7] https://github.com/XiaoxinHe/TAPE. For the GPT-3.5 and GPT-4, we simply call the API provided by OpenAI with default hyper-parameter settings. For the Llama2, we adopt the Llama-2-13b version and call their official implementation on https://huggingface.co. We empirically tried with some combinations of recommended important hyper-parameters, e.g., temperature and top_P, yet did not observe significant improvement. To realise the embedding-based information retrieval for LlmCorr, we adopt two capable embedding models ($f_{Emb}$) provided by OpenAI  [^8] https://platform.openai.com/docs/models/embeddings, e.g., text-embedding-3-large and text-embedding-3-small. In this work, we mainly adopt text-embedding-3-large for better empirical performance. We perform careful discussions about the influence of different variants in Section 5.2.
Appendix F Are LLMs Predictors?
@@
Figure 11: Zero-shot prompt templates for classification tasks.
Figure 12: Zero-shot prompt templates for regression tasks.
Figure 13: Few-shot prompt template for classification tasks. Multiple contextual knowledge can be included by expanding the template.
Figure 14: Few-shot prompt template for regression tasks. Multiple contextual knowledge can be included by expanding the template.
F.1 Predictor Prompt Engineering
Zero-shot Prompting. The first set of prompts (IP, IE) simply provides the LLM with molecule and task descriptions and asks it to generate the desired output with a desired format without any prior training or knowledge on the task, as illustrated in Figure 11 and Figure 12. The only guidance we provide to the LLM is instruction, which tells about a little background context. Particularly, IP only asks the LLM to provide predictions, while IE further asks for explanations, which may ask the LLM to clarify the thought process in explanation generation and provide helpful evidence to help users understand the given prediction. In addition, if we fill out the description of IP and IE, which derives IPD and IED prompts.
Few-shot Prompting. The second kind of prompt (FS) that we propose provides the LLM with a small number of examples of the task, along with the desired outputs Brown et al. (2020). The model then learns from these examples to perform the task on new inputs. This approach can be categorised as a simple in-context learning (ICL) technique, An example prompt template is shown in Figure 13 and Figure 14. FS-$k$ indicates $k$ contextual knowledge instances are included in the prompt. In this work, we do not discuss the FSD prompts since the generated descriptions have tons of tokens, which will easily go over the LLM’s input constraints.
We note there are also some popular recent ICL techniques, e.g., Chain-of-thought (CoT) Wei et al. (2022b), Tree-of-thought (ToT) Yao et al. (2023), Graph-of-thought (GoT) Besta et al. (2023) and Retrieval Augmented Generation (RaG) Lewis et al. (2020), which are theoretically available to support complicated tasks and include large knowledge context. However, our initial experiments showed that methods, e.g., CoT, ToT and GoT, perform much worse for molecule property prediction tasks due to the significant difficulties in designing proper chain thoughts without solid expertise. RaG implementations that we tested are unstable and slow with query, and they fall short of the relatively simpler FS’s performance. We argue the unqualified information retrieval system causes it, and we will discuss it in the future work discussion section.
F.2 Results - LLMs work as Predictors